### v3

@@ 5 Experiments
Qwen2-VL 100.0 21.0 0.0 23.3 38.3 25.3 0.0 29.3 97.0 23.3 0.0 17.3 39.0 30.0 0.0 15.3 61.0
RM Phi-3.5-vision 100.0 57.3 0.0 16.0 66.0 76.3 0.0 17.3 90.0 46.0 0.0 1.7 64.3 50.0 0.0 56.0 84.0
Qwen2-VL 100.0 19.0 0.0 15.7 25.3 43.0 0.0 60.3 84.0 5.3 0.0 5.7 8.3 58.0 0.0 60.3 80.7
PG Phi-3.5-vision 100.0 41.0 0.0 27.0 46.0 52.0 0.0 4.3 90.0 34.0 0.0 10.3 38.0 64.3 0.0 0.0 75.0
@@ 5.1 Experimental Setup
Victim Agents. Following prior studies Liu et al. (2024, 2023), we designed two representative digital agents: RecipeMaster (RM) and PoetryGenius (PG). RecipeMaster generates recipes based on uploaded ingredient images and user instructions, while PoetryGenius creates metrically sophisticated verses inspired by landscape photos and user preferences. Both agents incorporate external knowledge by accessing local documents and online resources. Their workflows involve multiple steps, including multimodal data processing, API calls, logical reasoning, and response generation. System prompts for the agents are provided in the supplementary material.
Tested VLMs. We adopt two state-of-the-art VLMs, Qwen2-VL Wang et al. (2024) and Phi-3.5-vision Abdin et al. (2024), as the core planners for the agents. Qwen2-VL is a novel vision-language model that demonstrates superior performance across various visual understanding benchmarks, which is well suited for integration into agent systems, such as mobile phones and physical robots Qwen Team (2024). Phi-3.5-vision, developed by Microsoft, achieves cutting-edge performance in image understanding while maintaining robust safety capabilities Ying et al. (2024). For both models, we set the max new tokens parameter to 1024.
Injected Tasks. Our evaluation leverages three public natual language processing datasets, which are unrelated to the agents’ predefined roles: CoEDIT Raheja et al. (2023) for text editing, SST2 Socher et al. (2013) for sentiment classification, and Code-to-Text Husain et al. (2020) for code explanation. For each dataset, 100 entries were randomly sampled to construct a injected instruction dataset. The results on Code-to-Text dataset are provided in supplementary material.
@@
Compared Attacks. We compare our method with two representative prompt injections. JIP Shayegani et al. (2024) implements visual modality-based prompt injection attacks on VLMs by embedding malicious instructions into the latent feature space of benign image inputs. UPI Liu et al. (2024) controls the agent’s output by appending adversarial suffixes to malicious instructions. Additionally, we formalize the direct instruction of agents to execute malicious instructions as the naive baseline method.
Evaluation Metrics. We evaluate the effectiveness of attacks using Attack Success Rate (ASR). A higher ASR indicates stronger adversarial control over the victim agents. We also use Performance under No Attack (PNA) to assess how well the agents perform their original tasks in the absence of attacks. A higher PNA indicates stronger capability of the agent in executing its designated tasks. Both metrics are calculated using the LLM-as-a-Judge approach Li et al. (2024), with corresponding prompt in supplementary material.
Implementation Details. We set the visual perturbation budget $\epsilon$ to 16 under the $\ell_{\infty}$ constraint, a value commonly used in visual adversarial examples Carlini et al. (2019), and the number of iterations to 200. We employ four types of multimodal vision encoders as surrogate models for visual attacks, including ViT-H-14, ViT-L-14, ViT-B-16, and ViT-SO400M-14-SigLIP-384. For textual attacks based on GCG algorithm, we set the top-k value to 256, batch size to 512, and the number of iterations to 100. Specifically, we leverage Llama-3.1-8B-Instruct as the surrogate large language model for optimizing adversarial textual commands. The LLM judge employs Qwen-Max Qwen Team (2024), a powerful large language model. To ensure statistical reliability, each experiment was repeated three times, and the average results are reported. All code is implemented in PyTorch, and all experiments are conducted on an NVIDIA A800-SXM4-80GB GPU cluster.
@@ 5.2 Main Results
Tab. 1 illustrates the overall evaluation results of our prompt injection method and the other attacks. We can identify: ❶ The consistent 100% performance under no attack condition directly demonstrates that the evaluated agents all maintain perfect capability and robustness when handling normal vision-language tasks within their roles. For instance, when processing standard recipe or poetry generation tasks within their designated functional roles, the agents consistently delivered correct outputs. ❷ The visual modality-based method JIP achieved a 0% ASR in all tested scenarios, demonstrating that merely embedding malicious instructions into benign image input fails to achieve injection attack effects on state-of-the-art multimodal agents. Even the malicious instructions are fully embedded into visual inputs, it’s still hard for multimodal agents to extract the embedded instructions from non-textual modalities. This underscores the inherent limitations of visual perturbations in executing effective prompt injection attacks against multimodal agents. ❸ The textual modality-based method UPI achieved marginal improvement over the naive methods in only six scenarios, with an average ASR increase of 5.4%. Its attack performance remains significantly inferior to our approach in most cases. This result empirically proves that only adversarial perturbations on textual modality are quite insufficient to effectively hijack multimodal agents to perform injected tasks.
Compared to existing baseline methods, our prompt injection method effectively injects malicious prompts into multimodal agents, achieving the highest ASRs across all evaluated scenarios. ❶ In local document attack scenarios, our method demonstrates an average performance gain up to 30% over all baseline approaches, with maximum ASR improvement reaching 71.7%. ❷ In online webpage attack scenarios, our method achieves an average improvement of 26.4% over baseline approaches, with maximum ASR enhancements reaching 34.0%. These results indicate that cross-modal adversarial manipulation targeting multiple inputs is required to strengthen the attack effectiveness of prompt injection against multimodal agents. ❸ Online webpage data with more complex information structures exhibited an average 10.9% ASR compared to the local document attack surface, as illustrated in Fig. 3. Only when redirecting the RecipeMaster (based on Qwen2-VL) to text-editing tasks did both attack surfaces achieve comparable effectiveness. This result validates the increased difficulty for multimodal agents in extracting injected instructions from raw external data.
@@ 5.3 Ablation Studies
Textual Enhancement. We also evaluate the effects of textual guidance enhancement. Four configurations are compared: (1) the full CrossInject  method, (2) CrossInject w/o the textual enhancement, (3) CrossInject with the malicious command optimized on victim agent’s real system prompt, and (4) CrossInject with the malicious command replaced by a random string of the same length. As shown in Tab. 4, removing the malicious command leads to an average drop of 24.8%, and maximum decrease of 47.0%. Replacing the optimized command with a random string not only fails to improve attack performance but also significantly degrades it in three cases. This suggests that non-semantic commands may disrupt the syntactic and semantic coherence of the prompt, distracting the agent from executing malicious task. To further validate the effectiveness of the inferred system prompt in textual enhancement, we optimized malicious command using the agent’s real system prompts and compared the ASR with our approaches. Malicious command optimized with real system prompt exhibits stronger attack than inferring by LLM. However, the improvement is quite marginal (+3.0% on average). Compared to it, our approach is more practical under black-box setting.
@@ 6 Physical World Agent Case Study
Vehicle Setup. We employ LIMO Robotics (2025), a commercial smart vehicle with vision language control capabilities (Qwen2-VL core) for physical world experiments. Given that a physical agent typically lacks external data interfaces, we simplify the process by directly delivering the injected malicious task to the vehicle as user command. The experimental environment consists of a manually constructed driving simulation track, with a stop sign positioned at the center of the track to restrict straight-through movement. The autonomous driving agent is forced to obey traffic regulations by its system prompt (i.e., You are a driving assistant that strictly adheres to traffic rules.), such as detouring when facing the stop sign.
Attack Implementation. We apply the visual latent alignment in the form of adversarial patches Liu et al. (2019) tailored to match the shape of the stop sign, ensuring its visual stealthiness. All parameters in textual guidance enhancement are consistent with those in Sec. 5. We define accelerate through road as the injected malicious instruction, which directly contradicts the agent’s system prompt.
Results Analysis. An illustration of the experimental case is shown in Fig. 5. We conduct the experiment under the same case for 10 times. When only injecting malicious textual instruction (naive attack in Sec. 5), the attack succeeded in preventing the vehicle from bypassing the stop sign in 4 instances. Under our CrossInject method, the ASR increased to 9 out of 10 attempts. This result demonstrates that CrossInject significantly disrupts real-world physical agents.
@@ 7 Countermeasures
To further assess the resilience of our cross-modal prompt injection attack, we evaluate representative defense strategies targeting both textual and visual modalities. These experiments are conducted on the Sentiment Analysis task, focusing on the local document input surface. For textual defense, we adopt the sandwich prompting Schulhoff (2024), a prevention-based method that inserts defensive prompts to reinforce the agent’s intended behavior and suppress adversarial instructions. For visual defense, due to the absence of existing methods specifically designed to counter visual injection, we adapt techniques originally developed for mitigating visual jailbreak attacks. Specifically, we apply Gaussian Blur Zhang et al. (2025) to the input image, using a $9\times 9$ kernel to reduce high-frequency details and potentially disrupt embedded adversarial patterns.
Textual Defense. We embed a defensive prompt within the external document input to reinforce the agent’s assigned role and mitigate attacker-specified instructions. As shown in Tab. 5, this strategy results in an average reduction of 6.7% in ASR, with a maximum reduction of 15.0%. These results suggest that contextual reinforcement through explicit role prompts can suppress conflicting adversarial commands by anchoring the agent to its original objective. The inserted defensive prompt likely improves the model’s resistance to task diversion by enhancing its awareness of the intended task context.
@@
Table 5: Defense results (ASR %) on our attack.
@@ 8 Conclusion and Future Work
Limitations. Though promising, we would like to explore the following aspects in the future: ❶ designing more effective defense strategies such as adversarial training Liu et al. (2023); Zhang et al. (2021); Liu et al. (2021) for multimodal setting in open environments, and ❷ extending attack applicability to real-world deployment scenarios with more heterogeneous environmental constraints.

### v4

@@ 5 Experiments
Qwen2-VL 100.0 21.0 0.0 25.0 38.3 25.3 0.0 28.0 97.0 23.3 0.0 15.0 39.0 30.0 0.0 60.0 61.0
RM Phi-3.5-vision 100.0 57.3 0.0 47.0 66.0 76.3 0.0 75.3 90.0 46.0 0.0 54.7 64.3 50.0 0.0 82.0 84.0
Qwen2-VL 100.0 19.0 0.0 20.0 25.3 43.0 0.0 62.0 84.0 5.3 0.0 6.0 8.3 58.0 0.0 73.0 80.7
PG Phi-3.5-vision 100.0 41.0 0.0 40.7 46.0 52.0 0.0 61.0 90.0 34.0 0.0 30.3 38.0 64.3 0.0 59.0 75.0
@@ 5.1 Experimental Setup
Victim Agents. Following prior studies Liu et al. (2024, 2023), we designed two representative digital chatbots for visual question answering Xie et al. (2024): RecipeMaster (RM) and PoetryGenius (PG). RecipeMaster generates recipes based on uploaded ingredient images and user instructions, while PoetryGenius creates metrically sophisticated verses inspired by landscape photos and user preferences. Both agents incorporate external knowledge by accessing local documents and online resources. Their workflows involve multiple steps, including multimodal data processing, API calls, logical reasoning, and response generation.
Tested VLMs. We adopt two state-of-the-art VLMs, Qwen2-VL Wang et al. (2024) with 2B parameters and Phi-3.5-vision Abdin et al. (2024), as the core planners for the agents. Qwen2-VL is a novel vision-language model that demonstrates superior performance across various visual understanding benchmarks, which is well suited for integration into agent systems, such as mobile phones and physical robots Qwen Team (2024). Phi-3.5-vision, developed by Microsoft, achieves cutting-edge performance in image understanding while maintaining robust safety capabilities Ying et al. (2024). For both models, we set the max new tokens parameter to 1024. Each agent can flexibly choose either Qwen2-VL or Phi-3.5-vision as its core planner.
Injected Tasks. Our evaluation leverages three public natural language processing datasets, which are unrelated to the agents’ predefined roles: CoEDIT Raheja et al. (2023) for text editing and SST2 Socher et al. (2013) for sentiment classification. For each dataset, 100 entries were randomly sampled to construct an injected instruction dataset.
@@
Compared Attacks. We compare our method with two representative prompt injections. JIP Shayegani et al. (2024) implements visual modality-based prompt injection attacks on VLMs by embedding malicious instructions into the latent feature space of benign image inputs. FB Liu et al. (2024) is a black-box query-based prompt injection attack that inserts textual delimiter into the input data to deceive agent to execute injected tasks. Additionally, we formalize the direct instruction of agents to execute malicious instructions as the Naive baseline method.
Evaluation Metrics. We evaluate the effectiveness of attacks using Attack Success Rate (ASR). A higher ASR indicates victim agents are more likely to respond to an attacker’s adversarial request. We also use Performance under No Attack (PNA) to assess how well the agents perform their original tasks in the absence of attacks. A higher PNA indicates stronger capability of the agent in executing its designated tasks. Both metrics are calculated using the LLM-as-a-Judge approach Li et al. (2024).
Implementation Details. We set the visual perturbation budget $\epsilon$ to 16 under the $\ell_{\infty}$ constraint, a value commonly used in visual adversarial examples Carlini et al. (2019), and the number of iterations to 200. We employ four types of multimodal vision encoders as surrogate models for visual attacks, including ViT-H-14, ViT-L-14, ViT-B-16, and ViT-SO400M-14-SigLIP-384. We use Stable-Diffusion-3.5-Large Rombach et al. (2022) to generate the target-aligned image. For textual attacks based on GCG algorithm, we set the top-k value to 256, batch size to 512, and the number of iterations to 100. Specifically, we leverage Llama-3.1-8B-Instruct as the surrogate large language model for optimizing adversarial textual commands. The LLM judge employs Qwen-Max Qwen Team (2024), a powerful large language model. To ensure statistical reliability, each experiment was repeated three times, and the average results are reported. All code is implemented in PyTorch, and all experiments are conducted on an NVIDIA A800-SXM4-80GB GPU cluster.
@@ 5.2 Main Results
Tab. 1 illustrates the overall evaluation results of our prompt injection method and the other attacks. We can identify: ❶ The consistent 100% performance under no attack condition directly demonstrates that the evaluated agents all maintain perfect capability and robustness when handling normal vision-language tasks within their roles. ❷ The visual modality-based method JIP achieved a 0% ASR in all tested scenarios, demonstrating that merely embedding malicious instructions into benign image input fails to achieve injection attack effects on state-of-the-art multimodal agents. Even the malicious instructions are fully embedded into visual inputs, it’s still hard for multimodal agents to extract the embedded instructions from non-textual modalities. This underscores the inherent limitations of visual perturbations in executing effective prompt injection attacks against multimodal agents. ❸ The textual modality-based method FB achieved limited improvement over the naive methods in ten cases, with an average ASR increase of 15.5%. Its attack performance remains significantly inferior to our approach in most cases. This result empirically proves that only adversarial perturbations on textual modality are quite insufficient to effectively hijack multimodal agents to perform injected tasks.
Compared to existing baseline methods, our prompt injection method effectively injects malicious prompts into multimodal agents, achieving the highest ASRs across all evaluated scenarios. ❶ In local document attack scenarios, our method demonstrates an average performance gain up to 32.7% over all baseline approaches, with maximum ASR improvement reaching 71.7%. ❷ In online webpage attack scenarios, our method achieves an average improvement of 27.5% over baseline approaches, with maximum ASR enhancements reaching 34.0%. These results indicate that cross-modal adversarial manipulation targeting multiple inputs is required to strengthen the attack effectiveness of prompt injection against multimodal agents. ❸ Online webpage data with more complex information structures exhibited an average 10.8% reduction on ASR compared to the local document attack surface, as illustrated in Fig. 3. Only when redirecting the RecipeMaster (based on Qwen2-VL) to text-editing tasks did both attack surfaces achieve comparable effectiveness. This result validates the increased difficulty for multimodal agents in extracting injected instructions from raw external data.
@@ 5.3 Ablation Studies
Textual Enhancement. We also evaluate the effects of textual guidance enhancement. Four configurations are compared: (1) the full CrossInject  method, (2) CrossInject w/o the textual enhancement, (3) CrossInject with the malicious command optimized on victim agent’s real system prompt, and (4) CrossInject with the malicious command replaced by a random string of the same length. As shown in Tab. 3, removing the malicious command leads to an average drop of 24.8%, and maximum decrease of 47.0%. Replacing the optimized command with a random string not only fails to improve attack performance but also significantly degrades it in three cases. This suggests that non-semantic commands may disrupt the syntactic and semantic coherence of the prompt, distracting the agent from executing malicious task. To further validate the effectiveness of the constructed system prompt in textual enhancement, we optimized malicious command using the agent’s real system prompts and compared the ASR with our approaches. Malicious command optimized with real system prompt exhibits stronger attack than constructed by LLM. However, the improvement is quite marginal (+3.0% on average). Compared to it, our approach is more practical under black-box setting.
[…]
Qwen2-VL 50.0 62.3 88.0 97.0
RM Phi-3.5-vision 81.0 64.7 92.3 90.0
Qwen2-VL 78.0 69.0 89.0 84.0
PG Phi-3.5-vision 76.0 43.0 91.7 90.0
@@ 6 Physical World Agent Case Study
Vehicle Setup. We employ LIMO Robotics (2025), a commercial smart vehicle with vision language control capabilities (Qwen2-VL core) for physical world experiments. Given that a physical agent typically lacks external data interfaces, we simplify the process by directly delivering the injected malicious task to the vehicle as user command. The experimental environment consists of a manually constructed driving simulation track, with a stop sign positioned at the center of the track to restrict straight-through movement. The autonomous driving agent is forced to obey traffic regulations by its system prompt, such as detouring when facing the stop sign.
Attack Implementation. We apply the visual latent alignment in the form of adversarial patches Liu et al. (2019) tailored to match the shape of the stop sign, ensuring its visual stealthiness. All parameters in textual guidance enhancement are consistent with those in Sec. 5.1. We define accelerate through road as the injected malicious instruction, which directly contradicts the agent’s system prompt.
Results Analysis. An illustration of the experimental case is shown in Fig. 5. We conduct the experiment under the same case for 10 times. When only injecting malicious textual instruction (Naive attack in Sec. 5.1), the attack succeeded in preventing the vehicle from bypassing the stop sign in 4 instances. Under our CrossInject method, the ASR increased to 9 out of 10 attempts. This result demonstrates that CrossInject significantly disrupts real-world physical agent.
@@ 7 Countermeasures
To further assess the resilience of our cross-modal prompt injection attack, we evaluate representative defense strategies targeting both textual and visual modalities. These experiments are conducted on the Sentiment Analysis task, focusing on the local document input surface. For textual defense, we adopt the sandwich prompting Schulhoff (2024), a prevention-based method that inserts defensive prompts to reinforce the agent’s intended behavior and suppress adversarial instructions. For visual defense, due to the absence of existing methods specifically designed to counter visual injection, we adapt techniques originally developed for mitigating visual jailbreak attacks. Specifically, we apply Gaussian Blur Zhang et al. (2025) to the input image, using a $\times 9$ 9 kernel to reduce high-frequency details and potentially disrupt embedded adversarial patterns.
Textual Defense. As shown in Tab. 5, textual-based defense results in an average reduction of 6.7% in ASR, with a maximum reduction of 15.0%. These results suggest that contextual reinforcement through explicit role prompts can suppress conflicting adversarial commands by anchoring the agent to its original objective. However, sandwich prompting cannot fully erase the attack efficacy. This may be attributed to the deceptive command adversarially optimized on constructed defensive prompt.
@@
Table 5: Defense results (ASR %) against our attack.
@@ 8 Conclusion and Future Work
Limitations. Though promising, we would like to explore the following aspects in the future: ❶ designing more effective defense strategies such as adversarial training Liu et al. (2023); Zhang et al. (2021); Liu et al. (2021) for multimodal setting in open environments, and ❷ extending attack applicability to real-world agents with more complex architecture.